OpenStudy (anonymous):

Given f(x)=x^2-1 and g(x)=x+1, find (fog)(x)?

OpenStudy (anonymous):

(fog)(x) = f(g(x)) = f(x+1)

OpenStudy (anonymous):

@shivraj u got what I did??????

OpenStudy (anonymous):

no sir

OpenStudy (anonymous):

that means you need to put x= x + 1 in f(x)

OpenStudy (anonymous):

Yep: (fog)=f(g(x))

OpenStudy (anonymous):

Now, substitute there g(x)

OpenStudy (anonymous):

suppose f(1) = 0 f(x) = x^2 - 1 f[g(x)] = (x+1)^2 - 1

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

you got it? @shivraj

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

if \[f(x)=x^2-1\ and \ g(x)=x+1, \\ \ then \ (fog)'(x)=\ (by \ product \ rule)\ u'v+u'v, \\where\ u=f(x); \ v=g(x) \\ \therefore\ (x^2-1)(1)+(2x)(x+1)=\\x^2-1+2x^2+2x=\\3x^2+2x-1=\\(3x-1)(x+1) \]

OpenStudy (anonymous):

thnks godfreysown